OpenStudy (devincoolage1001):

(x-7)/5=(x+3)/3 What is X

OpenStudy (legomyego180):

Any ideas how to start it?

OpenStudy (devincoolage1001):

I got 3x-21=5x+15, then 3x-6=5x. I am not sure if this is correct or where to go from there.

OpenStudy (adeevaabqari):

(x-7)/5 = (x+3)/3 3(x-7) = 5(x+3) 3x - 21 = 5x + 15 3x - 5x = 15 + 21 -2x = 36 x = 36/-2 x = -18

OpenStudy (devincoolage1001):

Thank you so much! This helped me a lot.

OpenStudy (adeevaabqari):

Yu're Welcome .. Glad to help you devin ^_^ ..